An aero plane is attempting to fly in a direction towards the east at a speed of 650kmh. A steady wind of speed of 90kmh is blowing from the north towards the south across the path of the aero plane. Determine the velocity of the plane with respect to the ground


1
Expert's answer
2022-01-31T07:34:01-0500


The resultant velocity


"V=\sqrt{650^2+90^2}=656\:\rm km/h""\theta=\tan^{-1} \frac{90}{650}=8^\circ\quad \rm S\;of\;E"